# Variations on five-dimensional sphere packings

### v2

@@ 5. Nine-dimensional kissing configurations
The kissing configuration contains $2882^{4}\cdot 18=$ 288 points that place arbitrary signs on the entries of the codewords, together with the $18$ permutations of (±$(\pm 2,0,\dots,0)$, for a total of $306$ points. A simple case analysis shows that no two distinct codewords overlap in more than two nonzero coordinates, which is exactly what is needed to obtain a valid kissing configuration.
@@
because $2=i^{2}$ and $i=(2+i)^{2}$, the elements $1$, $2$, $i$, and $2i$ are the nonzero squares, and we obtain the same code as before. The construction using $\mathbb{F}_{9}$ is more abstract, but it lets us see that the code is invariant under the affine semilinear group, which consists of the functions $x\mapsto ax+b$ and $x\mapsto ax^{3}+b$ with $a,b\in\mathbb{F}_{9}$ and $a\neq 0$. In fact, this group is the full automorphism group of the code. The automorphism group acts transitively on the codewords, since multiplying by a non-square interchanges the nonzero squares with the non-squares. By contrast, the subgroup generated by permuting the rows and columns and applying symmetries of the $3\times 3$ square does not act transitively.
@@
If we partition the kissing configuration into layers based on the first coordinate, we obtain layers of sizes $1$, $64$, $176$, $64$, and $1$ at heights $2$, $1$, 00, $-1$, and $-2$, respectively. To create a new kissing configuration, we replace the layer that has first coordinate $1$ with a modification. This mutation of an eight-dimensional layer is similar to how $L_{5}$, $Q_{5}$, and $R_{5}$ were constructed by modifying one four-dimensional layer of another five-dimensional kissing configuration, or to Figure 1.1.
Each point in the eight-dimensional layer to be modified has a $1$ in the first coordinate and three ±$\pm 1$’s in one of eight triples from among the remaining coordinates. One can check that there are exactly eight other triples that could be used instead while keeping the other layers unchanged, and we define the new kissing configuration by using them in this layer. [^4] Among the sixteen triples that are consistent with the central cross section, one can check that the original eight and their complement are the only sets of eight that avoid overlap in more than one coordinate, which would lead to overly large inner products. Another description is that in the modified layer, we swap coordinate $2$ with coordinate $3$ and coordinate $4$ with coordinate $7$. The modified points with first coordinate $1$ are therefore based on Table 5.2, rather than Table 5.1, by putting arbitrary signs on every coordinate except the first. [^5] Coordinates for both kissing configurations are available through DSpace@MIT at https://hdl.handle.net/1721.1/157699.
@@ Acknowledgements
We are grateful to Yoav Kallus for pointing out that the kissing configuration identified by Szöllősi in [20] had previously occurred as the kissing configuration of a packing discovered in [1].

### v3

@@ 5. Nine-dimensional kissing configurations
The kissing configuration contains $2^{4}\cdot 18=288$ points that place arbitrary signs on the entries of the codewords, together with the $18$ permutations of $(\pm 2,0,\dots,0)$, for a total of $306$ points. A simple case analysis shows that no two distinct codewords overlap in more than two nonzero coordinates, which is exactly what is needed to obtain a valid kissing configuration.
@@
because $2=i^{2}$ and $i=(2+i)^{2}$, the elements $1$, 22, $i$, and $2i$ are the nonzero squares, and we obtain the same code as before. The construction using $\mathbb{F}_{9}$ is more abstract, but it lets us see that the code is invariant under the affine semilinear group, which consists of the functions $x\mapsto ax+b$ and $x\mapsto ax^{3}+b$ with $a,b\in\mathbb{F}_{9}$ and $a\neq 0$. In fact, this group is the full automorphism group of the code. The automorphism group acts transitively on the codewords, since multiplying by a non-square interchanges the nonzero squares with the non-squares. By contrast, the subgroup generated by permuting the rows and columns and applying symmetries of the $3\times 3$ square does not act transitively.
@@
If we partition the kissing configuration into layers based on the first coordinate, we obtain layers of sizes $1$, $64$, $176$, $64$, and $1$ at heights $2$, 11, 0, 1-1, and 2-2, respectively. To create a new kissing configuration, we replace the layer that has first coordinate $1$ with a modification. This mutation of an eight-dimensional layer is similar to how $L_{5}$, $Q_{5}$, and $R_{5}$ were constructed by modifying one four-dimensional layer of another five-dimensional kissing configuration, or to Figure 1.1.
Each point in the eight-dimensional layer to be modified has a $1$ in the first coordinate and three $\pm 1$’s in one of eight triples from among the remaining coordinates. One can check that there are exactly eight other triples that could be used instead while keeping the other layers unchanged, and we define the new kissing configuration by using them in this layer. [^4] Among the sixteen triples that are consistent with the central cross section, one can check that the original eight and their complement are the only sets of eight that avoid overlap in more than one coordinate, which would lead to overly large inner products. Another description is that in the modified layer, we swap coordinate $2$ with coordinate $3$ and coordinate $4$ with coordinate $7$. The modified points with first coordinate $1$ are therefore based on Table 5.2, rather than Table 5.1, by putting arbitrary signs on every coordinate except the first. [^5] Coordinates for both kissing configurations are available through DSpace@MIT at https://hdl.handle.net/1721.1/157699.
@@ Acknowledgements
We are grateful to Yoav Kallus for pointing out that the kissing configuration identified by Szöllősi in [20] had previously occurred as the kissing configuration of a packing discovered in [1], and to the referees for helpful comments.